### v1

# Randomized block-Krylov subspace methods for low-rank approximation of matrix functions

Abstract
The randomized SVD is a method to compute an inexpensive, yet accurate, low-rank approximation of a matrix. The algorithm assumes access to the matrix through matrix-vector products (matvecs). Therefore, when we would like to apply the randomized SVD to a matrix function, $f(\bm{A})$, one needs to approximate matvecs with $f(\bm{A})$ using some other algorithm, which is typically treated as a black-box. Chen and Hallman (SIMAX 2023) argued that, in the common setting where matvecs with $f(\bm{A})$ are approximated using Krylov subspace methods (KSMs), more efficient low-rank approximation is possible if we open this black-box. They present an alternative approach that significantly outperforms the naive combination of KSMs with the randomized SVD, although the method lacked theoretical justification. In this work, we take a closer look at the method, and provide strong and intuitive error bounds that justify its excellent performance for low-rank approximation of matrix functions.
1 Introduction
Computing and approximating matrix functions is an important task in many application areas, such as differential equations [22, 24], network analysis [11, 13], and machine learning [9, 36]. For a symmetric matrix $\bm{A}\in\mathbb{R}^{n\times n}$ with eigenvalue decomposition
$\bm{A}=\bm{U}\bm{\Lambda}\bm{U}^{\textsf{{T}}},\quad\bm{\Lambda}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{n}),$
the matrix function is defined as
$f(\bm{A})=\bm{U}f(\bm{\Lambda})\bm{U}^{\textsf{{T}}},\quad f(\bm{\Lambda})=\operatorname{diag}(f(\lambda_{1}),\ldots,f(\lambda_{n})),$
where $f$ is a function defined on the eigenvalues of $\bm{A}$. Explicitly forming $f(\bm{A})$ generally costs $O(n^{3})$ operations, which becomes prohibitively expensive for large $n$.
However, instead of computing $f(\bm{A})$ exactly, in many cases we only need an approximation to $f(\bm{A})$, such as a low-rank approximation. For example, recent work has shown that low-rank approximations can be used to accurately estimate the trace or diagonal entries of $f(\bm{A})$ [3, 10, 26, 28, 31]. Furthermore, when $f(\bm{A})$ admits an accurate low-rank approximation $\bm{B}$, one can accurately and cheaply approximate matrix-vector products (matvecs) with $f(\bm{A})$ using the approximation
$\bm{B}\bm{Z}\approx f(\bm{A})\bm{Z}.$
Popular algorithms to compute cheap, yet accurate, low-rank approximations of matrices include the randomized SVD, randomized subspace iteration, and randomized block Krylov iteration. For an overview and analysis of these algorithms, we turn readers to [21, 29, 40]. These methods can be referred to as matrix-free, since they do not require explicit access to the matrix; they only require implicit access via matvecs. Furthermore, the dominant computational cost of these algorithms is usually the number of matvecs with the matrix that we want to approximate.
When the matrix we wish to approximate is a matrix-function, $f(\bm{A})$, we do not have access to exact matvecs. However, we can typically compute approximate matvecs through some black-box method, such as the (block) Lanczos method, or another Krylov subspace method [18, 20, 23, 4]. These methods are themselves matrix-free, only requiring matvecs with $\bm{A}$ to approximation matvecs with $f(\bm{A})$.
To summarize, the standard approach for efficiently computing a low-rank approximation to a matrix function, $f(\bm{A})$, combines two black-box methods: an outer-loop like the randomized SVD that requires matvecs with $f(\bm{A})$, and an inner-loop that computes each of these matvecs using multiple matvecs with $\bm{A}$ A broad theme of this paper and related work is that computational benefits can often be obtained by opening up these black-boxes. I.e., we can get away with fewer total matvecs with $\bm{A}$ to obtain a low-rank approximation to $f(\bm{A})$ if we shed the simple inner/outerloop approach. Such an approach was taken in [6] and [32, 33], where it is shown that algorithms like the randomized SVD and Nyström approximation for low-rank approximation, and Hutch++ for trace estimation, can benefit from such introspection.
Concretely, our work builds on a recent paper of Chen and Hallman [6], which develops so-called “Krylov-aware” variant of the Hutch++ and A-Hutch++ algorithms for trace estimation [28, 31] . The Hutch++ algorithm uses the randomized SVD to compute a low-rank approximation, which in turn is used reduce the variance of a stochastic estimator for the trace of $f(\bm{A})$ [8, 42, 43]. Chen and Hallman develop a more efficient method to compute this low-rank approximation for matrix functions. Numerical experiments suggest that their method is significantly more efficient than naively implementing the randomized SVD on $f(\bm{A})$ with approximate matvecs obtained in a black-box way from a Krylov subspace method. Our main contribution is to strengthen our theoretical understanding of the method from [6] by providing strong and intuitive error bounds that justify its excellent empirical performance.
1.1 Notation
Let $f:\mathbb{R}\mapsto\mathbb{R}$ be a function and let $\bm{A}$ be a symmetric matrix with eigendecomposition
$\begin{split}\bm{A}&=\bm{U}\bm{\Lambda}\bm{U}^{\textsf{{T}}}=\begin{bmatrix}\bm{U}_{k}&\bm{U}_{n\setminus k}\end{bmatrix}\begin{bmatrix}\bm{\Lambda}_{k}&\\ &\bm{\Lambda}_{n\setminus k}\end{bmatrix}\begin{bmatrix}\bm{U}_{k}^{\textsf{{T}}}\\ \bm{U}_{n\setminus k}^{\textsf{{T}}}\end{bmatrix},\\ \bm{\Lambda}_{k}&=\operatorname{diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{k}),\\ \bm{\Lambda}_{n\setminus k}&=\operatorname{diag}(\lambda_{k+1},\lambda_{k+2},\cdots,\lambda_{n}),\end{split}$ (1)
where the eigenvalues are ordered so that
$|f(\lambda_{1})|\geq|f(\lambda_{2})|\geq\ldots\geq|f(\lambda_{n})|.$
For a matrix $\bm{B}$ we denote by $\|\bm{B}\|_{\textsf{{F}}}$ and $\|\bm{B}\|_{2}$ the Frobenius norm and spectral norm, respectively. For a sketching matrix $\bm{\Omega}\in\mathbb{R}^{n\times\ell}$, where $\ell\geq k$, we define
$\bm{\Omega}_{k}:=\bm{U}_{k}^{\textsf{{T}}}\bm{\Omega},\quad\bm{\Omega}_{n\setminus k}:=\bm{U}_{n\setminus k}^{\textsf{{T}}}\bm{\Omega}$ (2)
For any matrix $\bm{B}\in\mathbb{R}^{m\times n}$, the submatrix consisting of rows $r_{1}$ through $r_{2}$ and columns $c_{1}$ through $c_{2}$ is denoted by $\bm{B}_{r_{1}:r_{2},c_{1}:c_{2}}$. We write $\bm{B}_{:,c_{1}:c_{2}}$ and $\bm{B}_{r_{1}:r_{2},:}$ to denote $\bm{B}_{1:m,c_{1}:c_{2}}$ and $\bm{B}_{r_{1}:r_{2},1:n}$ respectively. We write $\llbracket\bm{B}\rrbracket_{k}$ to denote the best rank $k$ approximation of $\bm{B}$ obtained by truncating the SVD. By $\mathbb{P}_{s}$ we denote the set of all polynomials whose degree is at most $s$. The Krylov subspace $\mathcal{K}_{s}(\bm{A},\bm{\Omega})$ is defined as
$\mathcal{K}_{s}(\bm{A},\bm{\Omega}):=\operatorname{range}\left(\begin{bmatrix}\bm{\Omega}&\bm{A}\bm{\Omega}&\cdots&\bm{A}^{s-1}\bm{\Omega}\end{bmatrix}\right).$
We denote $d_{s,\ell}:=\dim(\mathcal{K}_{s}(\bm{A},\bm{\Omega}))$. For a function $f$ defined on a set $I$ we write
$\|f\|_{L^{\infty}(I)}:=\sup\limits_{x\in I}|f(x)|.$
2 Krylov-aware low-rank approximation
We now describe and motivate the algorithm that we will analyse. Inspired by [6], we will call it Krylov-aware low-rank approximation. We begin with outlining the block-Lanczos algorithm to approximate matvecs with matrix functions. Next, we present how one would naively implement the randomized SVD on $f(\bm{A})$ using the block-Lanczos method to approximate matvecs. Finally, we present the alternative Krylov-aware algorithm and why this method allows us to gain efficiencies.
2.1 The block-Lanczos algorithm
Given an $n\times\ell$ matrix $\bm{\Omega}$, the block-Lanczos algorithm (Algorithm 1) can be used to iteratively obtain an orthonormal (graded) basis for $\mathcal{K}_{s}(\bm{A},\bm{\Omega})$. In particular, using $s$ block-matrix vector products with $\bm{A}$, the algorithm produces a basis $\bm{Q}_{s}$ and a block-tridiagonal matrix $\bm{T}_{s}$
$\bm{Q}_{s}=\begin{bmatrix}\bm{V}_{0}&\cdots&\bm{V}_{s-1}\end{bmatrix},\quad\bm{T}_{s}=\bm{Q}_{s}^{\textsf{{T}}}\bm{A}\bm{Q}_{s}=\operatorname{tridiag}\left(\begin{array}[]{c}\begin{array}[]{cccc}\bm{R}_{1}^{\textsf{{T}}}&\cdots&\bm{R}_{s-1}^{\textsf{{T}}}\end{array}\\ \begin{array}[]{ccccc}\bm{M}_{1}&\cdots&\cdots&\bm{M}_{s}\end{array}\\ \begin{array}[]{cccc}\bm{R}_{1}&\cdots&\bm{R}_{s-1}\end{array}\end{array}\right),$ (3)
where $\bm{R}_{0}$ is also output by the algorithm and is given by the relation $\bm{\Omega}=\bm{V}_{0}\bm{R}_{0}$.
Remark.
The block sizes of $\bm{Q}_{s}$ and $\bm{T}_{s}$ will never be larger than $\ell$, but may be smaller if the Krylov subspaces $\mathcal{K}_{i}(\bm{A},\bm{\Omega})$ obtained along the way have dimension less than $i\ell$. In particular, if $\mathcal{K}_{i}(\bm{A},\bm{\Omega})=\mathcal{K}_{i-1}(\bm{A},\bm{\Omega})$, then the blocks $\bm{V}_{j},\bm{R}_{j-1},\bm{M}_{j}$ will be empty for all $j\geq i-1$. The algorithm can terminate at this point, but for analysis it will be useful to imagine that the algorithm is continued and zero-width blocks are produced until it terminates at iteration $s$.
Algorithm 1 Block-Lanczos Algorithm
input: Symmetric $\bm{A}\in\mathbb{R}^{n\times n}$. Matrix $\bm{\Omega}\in\mathbb{R}^{n\times\ell}$. Number of iterations $s$.
output: Orthonormal basis $\bm{Q}_{s}$ for $\mathcal{K}_{s}(\bm{A},\bm{\Omega})$, and block tridiagonal $\bm{T}_{s}$.
1:Compute an orthonormal basis $\bm{V}_{0}$ for $\operatorname{range}(\bm{\Omega})$ and $\bm{R}_{0}=\bm{V}_{0}^{T}\bm{\Omega}$.
2:for $i=1,\ldots,s$ do
3:     $\bm{Y}=\bm{A}\bm{V}_{i-1}-\bm{V}_{i-2}\bm{R}_{i-1}^{\textsf{{T}}}$triangleright $\bm{Y}=\bm{A}\bm{V}_{i-1}$ if $i=1$
4:     $\bm{M}_{i}=\bm{V}_{i-1}^{\textsf{{T}}}\bm{Y}$
5:     $\bm{Y}=\bm{Y}-\bm{V}_{i-1}\bm{M}_{i}$
6:     $\bm{Y}=\bm{Y}-\sum_{j=0}^{i-1}\bm{V}_{j}\bm{V}_{j}^{\textsf{{T}}}\bm{Y}$triangleright reorthogonalize (repeat as needed)
7:     Compute an orthonormal basis $\bm{V}_{i}$ for $\operatorname{range}(\bm{Y})$ and $\bm{R}_{i}=\bm{V}_{i}^{T}\bm{Y}$.
8:end for
9:return $\bm{Q}_{s}$ and $\bm{T}_{s}$ as in Eq. 3
The block Lanczos algorithm can be used to approximate matvecs and quadratic forms with $f(\bm{A})$ using the approximations
$\bm{Q}_{s}f(\bm{T}_{s})_{:,1:\ell}\bm{R}_{0}$ $\approx f(\bm{A})\bm{\Omega},$ (4)
$\bm{R}_{0}^{\textsf{{T}}}f(\bm{T}_{s})_{1:\ell,1:\ell}\bm{R}_{0}$ $\approx\bm{\Omega}^{\textsf{{T}}}f(\bm{A})\bm{\Omega}.$ (5)
If $f$ is a low-degree polynomial, then the approximations Eqs. 4 and 5 are exact.111(4) and (5) are written out under the assumption that $\bm{\Omega}$ has rank $\ell$. If $\operatorname{rank}(\bm{\Omega})=r<\ell$, then the index set $1:\ell$ should be replaced with $1:r$ in both (4) and (5).
Lemma 2.1 ([6, Lemma 2.1]).
The approximation Eq. 4 is exact if $f\in\mathbb{P}_{s-1}$, and the approximation Eq. 5 is exact if $f\in\mathbb{P}_{2s-1}$.
It follows that Eqs. 4 and 5 are good approximations if $f$ is well approximated by polynomials. [^2] We note that for block-size $\ell>1$, the Krylov subspace is not equivalent to $\cup\{\operatorname{range}(p(\bm{A})\bm{\Omega}):p\in\mathbb{P}_{s-1}\}$, and bounds based on best approximation may be pessimistic due to this fact. In fact, deriving stronger bounds is an active area of research; see e.g. [5, 14, 15, 16, 17, 25]. In particular, one can obtain bounds in terms of the best polynomial approximation to $f$ on $[\lambda_{\min},\lambda_{\max}]$ [38, 30, 2].
2.2 The randomized SVD for matrix functions
The randomized SVD [21] is a simple and efficient algorithm to compute low-rank approximations of matrices that admit accurate low-rank approximations. The basic idea behind the randomized SVD is that if $\bm{\Omega}$ is a standard Gaussian random matrix, i.e. the entries in $\bm{\Omega}$ are independent identically distributed $N(0,1)$ random variables, then $\operatorname{range}(\bm{B}\bm{\Omega})$ should be reasonable approximation to the range of $\bm{B}$’s top singular vectors. Hence, projecting $\bm{B}$ onto $\operatorname{range}(\bm{B}\bm{\Omega})$ should yield a good approximation to $\bm{B}$. Algorithm 2 implements the randomized SVD on a symmetric matrix $\bm{B}$. The algorithm returns either the rank $\ell\geq k$ approximation $\bm{W}\bm{X}\bm{W}^{\textsf{{T}}}$ or the rank $k$ approximation $\bm{W}\llbracket\bm{X}\rrbracket_{k}\bm{W}^{\textsf{{T}}}$, depending on the needs of the user.
Algorithm 2 Randomized SVD
input: Symmetric $\bm{B}\in\mathbb{R}^{n\times n}$. Rank $k$. Oversampling parameter $\ell-k$.
output: Low-rank approximation to $\bm{B}$: $\bm{W}\bm{X}\bm{W}^{\textsf{{T}}}$ or $\bm{W}\llbracket\bm{X}\rrbracket_{k}\bm{W}^{\textsf{{T}}}$
1:Sample a standard Gaussian $n\times\ell$ matrix $\bm{\Omega}$.
2:Compute $\bm{K}=\bm{B}\bm{\Omega}$.
3:Compute an orthonormal basis $\bm{W}$ for $\operatorname{range}(\bm{K})$.
4:Compute $\bm{X}=\bm{W}^{\textsf{{T}}}\bm{B}\bm{W}$.
5:return $\bm{W}\bm{X}\bm{W}^{\textsf{{T}}}$ or $\bm{W}\llbracket\bm{X}\rrbracket_{k}\bm{W}^{\textsf{{T}}}$
Typically, the dominant cost of Algorithm 2 is that computation of matvecs $\bm{B}$. We require $2\ell$ such matvecs: $\ell$ in 2 with $\bm{\Omega}$ and $\ell$ in 4 with $\bm{W}$. When Algorithm 2 is applied to a matrix function $\bm{B}=f(\bm{A})$ these matvecs cannot be performed exactly, but need to be approximated using, for example, the block Lanczos method discussed in the previous section. Algorithm 3 implements the randomized SVD applied to $f(\bm{A})$ with approximate matvecs using the block Lanczos method. The cost is now $(s+r)\ell$ matvecs with $\bm{A}$, where $s$ and $r$ should be set sufficiently large so that the approximations Eqs. 4 and 5 are accurate.
Algorithm 3 Randomized SVD on a matrix function $f(\bm{A})$
input: Symmetric matrix $\bm{A}\in\mathbb{R}^{n\times n}$. Rank $k$. Oversampling parameter $\ell-k$. Matrix function $f:\mathbb{R}\to\mathbb{R}$. Accuracy parameters $s$ and $r$.
output: Low-rank approximation to $f(\bm{A})$: $\bm{W}\bm{X}\bm{W}^{\textsf{{T}}}$ or $\bm{W}\llbracket\bm{X}\rrbracket_{k}\bm{W}^{\textsf{{T}}}$
1:Sample a standard Gaussian $n\times\ell$ matrix $\bm{\Omega}$.
2:Run Algorithm 1 for $s$ iterations to obtain an orthonormal basis $\bm{Q}_{s}$ for $\mathcal{K}_{s}(\bm{A},\bm{\Omega})$, a block tridiagonal matrix $\bm{T}_{s}$ and an upper triangular matrix $\bm{R}_{0}$.
3:Compute the approximation $\bm{K}=\bm{Q}_{s}f(\bm{T}_{s})_{:,1:\ell}\bm{R}_{0}\approx f(\bm{A})\bm{\Omega}$.
4:Compute an orthonormal basis $\bm{W}$ for $\operatorname{range}(\bm{K})$.
5:Run Algorithm 1 for $r$ iterations with starting block $\bm{W}$ to obtain the block tridiagonal matrix $\widetilde{\bm{T}}_{r}$.
6:Compute the approximation $\bm{X}=f(\widetilde{\bm{T}}_{r})_{1:\ell,1:\ell}\approx\bm{W}^{\textsf{{T}}}f(\bm{A})\bm{W}$.
7:return $\bm{W}\bm{X}\bm{W}^{\textsf{{T}}}$ or $\bm{W}\llbracket\bm{X}\rrbracket_{k}\bm{W}^{\textsf{{T}}}$
2.3 Krylov-aware low-rank approximation
A key observation in [6] was that $\operatorname{range}(\bm{W})\subseteq\operatorname{range}(\bm{Q}_{s})$, where $\bm{W}$ and $\bm{Q}_{s}$ are as in Algorithm 3. Therefore, by [33, Lemma 3.3] one has
$\|f(\bm{A})-\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}$ $\leq\|f(\bm{A})-\bm{W}\bm{W}^{\textsf{{T}}}f(\bm{A})\bm{W}\bm{W}^{\textsf{{T}}}\|_{\textsf{{F}}},$
$\|f(\bm{A})-\bm{Q}_{s}\llbracket\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}$ $\leq\|f(\bm{A})-\bm{W}\llbracket\bm{W}^{\textsf{{T}}}f(\bm{A})\bm{W}\rrbracket_{k}\bm{W}^{\textsf{{T}}}\|_{\textsf{{F}}}.$
Hence, assuming that the quadratic form $\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}$ can be computed accurately, the naive implementation of the randomized SVD outlined in Algorithm 3 will yield a worse error than using $\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}$ as an approximation to $f(\bm{A})$.
Since $\bm{Q}_{s}$ could have as many as $s\ell$ columns, an apparent downside to this approach is that approximating $f(\bm{A})\bm{Q}_{s}$ might require $rs\ell$ matvecs with $\bm{A}$ if we use $r$ iterations of the block Lanczos method (Algorithm 1) to approximately compute each matvec with $f(\bm{A}$. The key observation which allows Krylov-aware algorithms to be implemented efficiently is the following:
Lemma 2.2 ([6, Section 3]).
Suppose that $\bm{Q}_{s}$ is the output of Algorithm 1 with starting block $\bm{\Omega}$ and $s$ iterations. Then, running $r+1$ iterations of Algorithm 1 with starting block $\bm{Q}_{s}$ yields the same output as running $s+r$ iterations of Algorithm 1 with starting block $\bm{\Omega}$.
The insight behind Lemma 2.2 is that, since $\bm{Q}_{s}$ is a basis for the Krylov subspace $\mathcal{K}_{s}(\bm{A},\bm{\Omega})$, we have $\mathcal{K}_{r+1}(\bm{A},\bm{Q}_{s})=\mathcal{K}_{s+r}(\bm{A},\bm{\Omega})$. This fact can be verified by explicitly writing out $\mathcal{K}_{r+1}(\bm{A},\bm{Q}_{s})$:
$\mathcal{K}_{r+1}(\bm{A},\bm{Q}_{s})$ $=\operatorname{range}\big{(}\big{[}\bm{Q}_{s}\,\bm{A}\bm{Q}_{s}\,\cdots\,\bm{A}^{r}\bm{Q}_{s}\big{]}\big{)}$
$=\operatorname{range}\big{(}\big{[}\bm{\Omega}\,\hskip 5.0pt\bm{A}\bm{\Omega}\,\hskip 5.0pt\cdots\hskip 5.0pt\,\bm{A}^{r}\bm{\Omega}$
$\hskip 62.59596pt\bm{A}\bm{\Omega}\,\hskip 5.0pt\bm{A}^{2}\bm{\Omega}\,\cdots\,\bm{A}^{r+1}\bm{\Omega}$
[…]
$\hskip 99.58464pt\bm{A}^{r}\bm{\Omega}\,\hskip 5.0pt\bm{A}^{r+1}\bm{\Omega}\,\hskip 5.0pt\cdots\,\bm{A}^{s+r-1}\bm{\Omega}\big{]}\big{)}=\mathcal{K}_{s+r}(\bm{A},\bm{\Omega}).$
In addition to the work of Chen and Hallman, this observation was recently used to analyze “small-block” randomized Krylov methods for low-rank approximation [27] and a randomized variant of the block conjugate gradient algorithm [7].
Notably, Lemma 2.2 enables us to approximate $\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}$ with just $r\ell$ additional matvecs with $\bm{A}$, even though $\bm{Q}_{s}$ has $s\ell$ columns! Hence, approximately computing $\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}$ is essentially no more expensive, in terms of the number of matvecs with $\bm{A}$, than approximating $\bm{W}^{\textsf{{T}}}f(\bm{A})\bm{W}$, as done in Algorithm 3 6. As a consequence of Lemma 2.2, we have the following result.
Lemma 2.3.
Let $\lambda_{\max}$ and $\lambda_{\min}$ denote the largest and smallest eigenvalue of $\bm{A}$. Let $q=s+r$ and let $\bm{T}_{q}$ and $\bm{Q}_{s}$ be computed using Algorithm 1. Then,
$\|\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}-f(\bm{T}_{q})_{1:d_{s,\ell},1:d_{s,\ell}}\|_{\textsf{{F}}}\leq 2\sqrt{\ell s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|f(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}.$
Proof.
By Lemmas 2.1 and 2.2 we know that for any polynomial $p\in\mathbb{P}_{2r+1}$ we have $\bm{Q}_{s}^{\textsf{{T}}}p(\bm{A})\bm{Q}_{s}=p(\bm{T}_{q})_{1:d_{s,\ell},1:d_{s,\ell}}$. Therefore, since $\|\bm{Q}_{s}\|_{\textsf{{F}}}\leq\sqrt{\ell s}$ and $\|\bm{Q}_{s}\|_{2}\leq 1$ we have
$\|\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}-f(\bm{T}_{q})_{1:d_{s,\ell},1:d_{s,\ell}}\|_{\textsf{{F}}}$
$=\|\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}-p(\bm{T}_{q})_{1:d_{s,\ell},1:d_{s,\ell}}+p(\bm{T}_{q})_{1:d_{s,\ell},1:d_{s,\ell}}-f(\bm{T}_{q})_{1:d_{s,\ell},1:d_{s,\ell}}\|_{\textsf{{F}}}$
$\leq\|\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}-\bm{Q}_{s}^{\textsf{{T}}}p(\bm{A})\bm{Q}_{s}\|_{\textsf{{F}}}+\|(p(\bm{T}_{q})-f(\bm{T}_{q}))_{1:d_{s,\ell},1:d_{s,\ell}}\|_{\textsf{{F}}}$
$\leq\sqrt{\ell s}\left(\|f(\bm{A})-p(\bm{A})\|_{2}+\|p(\bm{T}_{q})-f(\bm{T}_{q})\|_{2}\right)$
$\leq 2\sqrt{\ell s}\|f(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])},$
where the last inequality is due to the fact that the spectrum of $\bm{T}_{q}$ is contained in $[\lambda_{\min},\lambda_{\max}]$. Optimizing over $p\in\mathbb{P}_{2r+1}$ gives the result. ∎
We can now present the Krylov-aware low-rank approximation algorithm; see Algorithm 4. The total number of matvecs with $\bm{A}$ is $(s+r)\ell$, the same as Algorithm 3. However, as noted above, Algorithm 4 (approximately) projects $f(\bm{A})$ onto a higher dimensional subbspace, ideally obtaining a better approximation.
Algorithm 4 Krylov-aware low-rank approximation
input: Symmetric $\bm{A}\in\mathbb{R}^{n\times n}$. Rank $k$. Oversampling parameter $\ell-k$. Matrix function $f:\mathbb{R}\to\mathbb{R}$. Number of iterations $q=s+r$.
output: Low-rank approximation to $f(\bm{A})$: $\bm{Q}_{s}\bm{X}\bm{Q}_{s}^{\textsf{{T}}}$ or $\bm{Q}_{s}\llbracket\bm{X}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}$
1:Sample a standard Gaussian $n\times\ell$ matrix $\bm{\Omega}$.
2:Run Algorithm 1 for $q=s+r$ iterations to obtain an orthonormal basis $\bm{Q}_{s}$ for $\mathcal{K}_{s}(\bm{A},\bm{\Omega})$ and a block tridiagonal matrix $\bm{T}_{q}$.
3:Compute $\bm{X}=f(\bm{T}_{q})_{1:d_{s,\ell},1:d_{s,\ell}}$. $\triangleright$ $\approx\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}$
4:return $\textsf{{ALG}}(s,r;f)=\bm{Q}_{s}\bm{X}\bm{Q}_{s}^{\textsf{{T}}}$ or $\textsf{{ALG}}_{k}(s,r;f)=\bm{Q}_{s}\llbracket\bm{X}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}$.
We conclude by noting that the function $f$ in Algorithm 4 does not need to be fixed; one can compute a low-rank approximation for many different functions $f$ at minimal additional cost.
3 Error bounds
In this section, we establish error bounds for Algorithm 4. We break the analysis into two parts. First, in Section 3.1, we derive general error bounds for approximations to $f(\bm{A})$ when projections $\bm{Q}\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\bm{Q}^{\textsf{{T}}}$ cannot be computed exactly. In Section 3.2 we provide structural bounds for the errors $\|f(\bm{A})-\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}$ and $\|f(\bm{A})-\bm{Q}_{s}\llbracket\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}$ that hold with probability $1$, and in Section 3.3 we derive the corresponding probabilistic bounds. Next, in Section 3.4 we combine the results from Sections 3.1, 3.2 and 3.3 to derive end-to-end error bounds for Algorithm 4, which involves approximate projection onto $\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}$.
3.1 Error bounds for inexact projections
In this section we will derive error bounds for $\|f(\bm{A})-\bm{Q}\bm{X}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}$ and $\|f(\bm{A})-\bm{Q}\llbracket\bm{X}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}$ where $\bm{Q}$ is any orthonormal basis and $\bm{X}$ is any matrix. By [33, Lemma 3.3] we know that the optimal choice of $\bm{X}$ is $\bm{X}=\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}$. However, since $\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}$ can only be computed approximately, we need to show that the errors $\|f(\bm{A})-\bm{Q}\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\bm{Q}\|_{\textsf{{F}}}$ and $\|f(\bm{A})-\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\rrbracket_{k}\bm{Q}\|_{\textsf{{F}}}$ are robust against perturbations in $\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}$. Theorem 3.1 provides such a result.
Theorem 3.1.
Given an orthonormal basis $\bm{Q}$ and an approximation $\bm{X}$ to $\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}$. Then,
$\|f(\bm{A})-\bm{Q}\bm{X}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}^{2}=\|f(\bm{A})-\bm{Q}\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}^{2}+\|\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}-\bm{X}\|_{\textsf{{F}}}^{2},$ (6)
and
$\|f(\bm{A})-\bm{Q}\llbracket\bm{X}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}\leq\|f(\bm{A})-\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}+2\|\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}-\bm{X}\|_{\textsf{{F}}}.$ (7)
Proof.
(6) is immediate due to the Pythagorean theorem.
We now proceed with proving (7) using a similar argument to [41, Proof of Theorem 5.1]. Define $\bm{C}=f(\bm{A})-\bm{Q}\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\bm{Q}^{\textsf{{T}}}+\bm{Q}\bm{X}\bm{Q}^{\textsf{{T}}}$. Note that $\|\bm{C}-f(\bm{A})\|_{\textsf{{F}}}=\|\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}-\bm{X}\|_{\textsf{{F}}}$ and $\bm{Q}^{\textsf{{T}}}\bm{C}\bm{Q}=\bm{X}$. Hence,
$\|f(\bm{A})-\bm{Q}\llbracket\bm{X}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}$ $=\|f(\bm{A})-\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}\bm{C}\bm{Q}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}$
$\leq\|f(\bm{A})-\bm{C}\|_{\textsf{{F}}}+\|\bm{C}-\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}\bm{C}\bm{Q}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}$
$=\|\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}-\bm{X}\|_{\textsf{{F}}}+\|\bm{C}-\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}\bm{C}\bm{Q}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}.$ (8)
Since $\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}\bm{C}\bm{Q}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}$ is the best rank $k$ approximation whose range is contained in $\operatorname{range}(\bm{Q})$ [33, Lemma 3.3] (a similar result is given in [19, Theorem 3.5]), we have
$\|\bm{C}-\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}\bm{C}\bm{Q}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}\leq$ $\|\bm{C}-\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}}$
$\leq$ $\|\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}-\bm{X}\|_{\textsf{{F}}}+\|f(\bm{A})-\bm{Q}\llbracket\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}\rrbracket_{k}\bm{Q}^{\textsf{{T}}}\|_{\textsf{{F}}},$ (9)
where the second inequality is due to the fact that $\|f(\bm{A})-\bm{C}\|_{\textsf{{F}}}=\|\bm{Q}^{\textsf{{T}}}f(\bm{A})\bm{Q}-\bm{X}\|_{\textsf{{F}}}$. Combining (8) and (9) yields the desired inequality. ∎
A corollary of Theorem 3.1 is that the error of the approximation from Algorithm 4 will always be bounded from above by the error of the approximation from Algorithm 3, up to a polynomial approximation error of $f$:
Corollary 3.2.
Let $\lambda_{\max}$ and $\lambda_{\min}$ be the largest and smallest eigenvalues of $\bm{A}$. Let $\bm{Q}_{s}$ and $\bm{X}$ be the output from Algorithm 4 and let $\bm{W}$ and $\widetilde{\bm{X}}$ be the output from Algorithm 3 with the same input parameters and the same sketch matrix $\bm{\Omega}$. Then,
$\|f(\bm{A})-\bm{Q}_{s}\llbracket\bm{X}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}$
$\hskip 20.00003pt\leq\|f(\bm{A})-\bm{W}\llbracket\widetilde{\bm{X}}\rrbracket_{k}\bm{W}^{\textsf{{T}}}\|_{\textsf{{F}}}+4\sqrt{\ell s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|f(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}.$
Proof.
Since $\operatorname{range}(\bm{Q}_{s})\subseteq\operatorname{range}(\bm{W})$ we have
$\|f(\bm{A})-\bm{Q}_{s}\llbracket\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}$
$\hskip 70.0001pt\leq\|f(\bm{A})-\bm{W}\llbracket\bm{W}^{\textsf{{T}}}f(\bm{A})\bm{W}\rrbracket_{k}\bm{W}^{\textsf{{T}}}\|_{\textsf{{F}}}$
$\hskip 140.00021pt\leq\|f(\bm{A})-\bm{W}\llbracket\widetilde{\bm{X}}\rrbracket_{k}\bm{W}^{\textsf{{T}}}\|_{\textsf{{F}}},$
where we used that $\bm{W}\llbracket\bm{W}^{\textsf{{T}}}f(\bm{A})\bm{W}\rrbracket_{k}\bm{W}^{\textsf{{T}}}$ is the best rank $k$ approximation to $f(\bm{A})$ whose range and co-range is contained in $\operatorname{range}(\bm{W})$ [33, Lemma 3.3]. Hence, by Theorem 3.1 we have
$\|f(\bm{A})-\bm{Q}_{s}\llbracket\bm{X}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}$ $\leq\|f(\bm{A})-\bm{Q}_{s}\llbracket\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}+2\|\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}-\bm{X}\|_{\textsf{{F}}}$
$\leq\|f(\bm{A})-\bm{W}\llbracket\widetilde{\bm{X}}\rrbracket_{k}\bm{W}^{\textsf{{T}}}\|_{\textsf{{F}}}+2\|\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}-\bm{X}\|_{\textsf{{F}}}.$
Applying Lemma 2.3 yields the desired result. ∎
Corollary 3.2 already shows why we expect the Krylov-aware approach in Algorithm 4 to outperform a naive combination of Krylov subspace methods and the randomized SVD in Algorithm 3. In fact, we might expect a major improvement, since $\operatorname{range}(\bm{Q}_{s})$ is a significantly larger subspace than $\operatorname{range}(\bm{W})$. In the subsequent sections we will derive stronger bounds for the approximation returned by Algorithm 4 to better justify this intuition.
3.2 Structural bounds
We next derive structural bounds for $\|f(\bm{A})-\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}$ and $\|f(\bm{A})-\bm{Q}_{s}\llbracket\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}$ that is true for any sketch matrix $\bm{\Omega}$ as long as $\bm{\Omega}_{k}$ defined in (2) has rank $k$. These bounds will allow us to obtain probabilistic bounds on the error of approximating $f(\bm{A})$, at least under the assumption that $\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}$ and $\bm{Q}_{s}\llbracket\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}$ are computed exactly. As a reminder, we will remove this assumption in Section 3.4 using the perturbation bounds from Section 3.1.
To state our bounds, we introduce the quantity
$\mathcal{E}_{\bm{\Omega}}(s;f)=\min\limits_{p\in\mathbb{P}_{s-1}}\left[\|p(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}\left|\frac{f(\lambda_{i})}{p(\lambda_{i})}\right|^{2}\right],$ (10)
which quantifies the extent to which an $(s-1)$ degree polynomial can be large (relative to $f$) on the eigenvalues $\lambda_{1},\ldots,\lambda_{k}$ and small on the remaining eigenvalues. Recall that we order $\bm{A}$’s eigenvalues with respect to $f$, so $|f(\lambda_{1})|\geq|f(\lambda_{1})|\geq\ldots\geq|f(\lambda_{n})|$.
Lemma 3.3.
Consider $\bm{A}\in\mathbb{R}^{n\times n}$ as defined in (1). Assuming $\bm{\Omega}_{k}$ in (2) has rank $k$, for all functions $f:\mathbb{R}\to\mathbb{R}$, we have
$\begin{split}&\|f(\bm{A})-\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}^{2}\\ &\hskip 70.0001pt\leq\|f(\bm{A})-\bm{Q}_{s}\llbracket\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}^{2}\\ &\hskip 140.00021pt\leq\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+5\mathcal{E}_{\bm{\Omega}}(s;f).\end{split}$ (11)
From this bound, we can further see why Algorithm 4 should be preferred over directly applying the randomized SVD to $f(\bm{A})$. As an extreme case, consider when $f$ is a polynomial of degree at most $s-1$. Then the standard randomized SVD bound [21, Theorem 9.1] essentially replaces $p$ with $f$ in $\mathcal{E}_{\bm{\Omega}}(s;f)$ in (11):
$\|f(\bm{A})-\bm{W}\llbracket\bm{W}^{T}f(\bm{A})\bm{W}\rrbracket_{k}\bm{W}\|_{\textsf{{F}}}^{2}\leq\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+5\|f(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2},$
where $\bm{W}$ is an orthonormal basis for $\operatorname{range}(f(\bm{A})\bm{\Omega})$. Since $\mathcal{E}_{\bm{\Omega}}(s;f)$ minimizes over all degree $s-1$ polynomials, it is always smaller than $\|f(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2}$. So, the approximation returned by Algorithm 4 is expected to be more accurate compared to the randomized SVD. When $f$ is not a polynomial, the error of the randomized SVD roughly corresponds to plugging a good polynomial approximation for $f$ into Eq. 10, but again there might be a better choice to minimize $\mathcal{E}_{\bm{\Omega}}(s;f)$.
Indeed, in many cases, we find that the improvement obtained from effectively optimizing over all possible degree $s-1$ polynomials is significant – there is often a much better choice of polynomial than a direct approximation to $f$ for Eq. 10. This is reflected in our experiments (Section 4) and we also provide a concrete analysis involving the matrix exponential in Section 3.5.
Proof of Lemma 3.3.
The first inequality is due to the fact that $\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}$ is the nearest matrix to $f(\bm{A})$ in the Frobenius norm whose range and co-range is contained in $\operatorname{range}(\bm{Q}_{s})$ [33, Lemma 3.3]. We proceed with proving the second inequality.
To do so, it of course suffices to prove the inequality where $\mathcal{E}_{\bm{\Omega}}(s;f)$ is replaced with $\|p(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2}\cdot\max_{i=1,\ldots,k}\left|{f(\lambda_{i})}/{p(\lambda_{i})}\right|^{2}$ for any choice of $p\in\mathbb{P}_{s-1}$. Note that if we choose $p$ so that $p(\lambda_{i})=0$ for some $i=1,\ldots,k$ then the right hand side of (11) is infinite and the bound trivially holds. Hence, we may assume that $p(\lambda_{i})\neq 0$ for $i=1,\ldots,k$. Consequently, $p(\bm{\Lambda}_{k})$ is non-singular. Define $\bm{Z}=p(\bm{A})\bm{\Omega}\bm{\Omega}_{k}^{\dagger}p(\bm{\Lambda}_{k})^{-1}$ and let $\widetilde{\bm{P}}$ be the orthogonal projector onto $\operatorname{range}(\bm{Z})\subseteq\operatorname{range}(\bm{Q}_{s})$. Note that $\operatorname{rank}(\bm{Z})\leq k$ and $\bm{Q}_{s}\llbracket\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}$ is the best rank $k$ approximation to $f(\bm{A})$ whose range and co-range is contained in $\operatorname{range}(\bm{Q}_{s})$ [33, Lemma 3.3]. Hence,
$\|f(\bm{A})-\bm{Q}_{s}\llbracket\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}^{2}\leq\|f(\bm{A})-\widetilde{\bm{P}}f(\bm{A})\widetilde{\bm{P}}\|_{\textsf{{F}}}^{2}.$
Now define $\widehat{\bm{P}}=\bm{U}^{\textsf{{T}}}\widetilde{\bm{P}}\bm{U}$, which is the orthogonal projector onto $\operatorname{range}(\bm{U}^{\textsf{{T}}}\bm{Z})$. By the unitary invariance of the Frobenius norm we have
$\|f(\bm{A})-\widetilde{\bm{P}}f(\bm{A})\widetilde{\bm{P}}\|_{\textsf{{F}}}^{2}=\|f(\bm{\Lambda})-\widehat{\bm{P}}f(\bm{\Lambda})\widehat{\bm{P}}\|_{\textsf{{F}}}^{2}.$
Furthermore, by Pythagorean theorem,
$\|f(\bm{\Lambda})-\widehat{\bm{P}}f(\bm{\Lambda})\widehat{\bm{P}}\|_{\textsf{{F}}}^{2}=\|(\bm{I}-\widehat{\bm{P}})f(\bm{\Lambda})\|_{\textsf{{F}}}^{2}+\|\widehat{\bm{P}}f(\bm{\Lambda})(\bm{I}-\widehat{\bm{P}})\|_{\textsf{{F}}}^{2}.$ (12)
We are going to bound the two terms on the right hand side of (12) separately. Our analysis for the first is similar to the proof of [21, Theorem 9.1].
Note that since $\operatorname{rank}(\bm{\Omega}_{k})=k$ we have $\bm{\Omega}_{k}\bm{\Omega}_{k}^{\dagger}=\bm{I}$. Hence,
$\bm{U}^{\textsf{{T}}}\bm{Z}=\bm{U}^{\textsf{{T}}}p(\bm{A})\bm{\Omega}\bm{\Omega}_{k}^{\dagger}p(\bm{\Lambda}_{k})^{-1}=\begin{bmatrix}\bm{I}\\ p(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}p(\bm{\Lambda}_{k})^{-1}\end{bmatrix}=:\begin{bmatrix}\bm{I}\\ \bm{F}\end{bmatrix}.$
Hence,
$\bm{I}-\widehat{\bm{P}}$ $=\begin{bmatrix}\bm{I}-(\bm{I}+\bm{F}^{\textsf{{T}}}\bm{F})^{-1}&-(\bm{I}+\bm{F}^{\textsf{{T}}}\bm{F})^{-1}\bm{F}^{\textsf{{T}}}\\ -\bm{F}(\bm{I}+\bm{F}^{\textsf{{T}}}\bm{F})^{-1}&\bm{I}-\bm{F}(\bm{I}+\bm{F}^{\textsf{{T}}}\bm{F})^{-1}\bm{F}^{\textsf{{T}}}\end{bmatrix}$
$\preceq\begin{bmatrix}\bm{F}^{\textsf{{T}}}\bm{F}&-(\bm{I}+\bm{F}^{\textsf{{T}}}\bm{F})^{-1}\bm{F}^{\textsf{{T}}}\\ -\bm{F}(\bm{I}+\bm{F}^{\textsf{{T}}}\bm{F})^{-1}&\bm{I}\end{bmatrix},$
where the inequality is due to [21, Proposition 8.2]. Consequently,
$\|(\bm{I}-\widehat{\bm{P}})f(\bm{\Lambda})\|_{\textsf{{F}}}^{2}$ $=\operatorname{tr}(f(\bm{\Lambda})(\bm{I}-\widehat{\bm{P}})f(\bm{\Lambda}))$
$\leq\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+\|\bm{F}f(\bm{\Lambda}_{k})\|_{\textsf{{F}}}^{2}$
$\leq\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+\|p(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2}\|p(\bm{\Lambda}_{k})^{-1}f(\bm{\Lambda}_{k})\|_{2}^{2}$
$=\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+\|p(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}\left|\frac{f(\lambda_{i})}{p(\lambda_{i})}\right|^{2}.$ (13)
We proceed with bounding the second term in (12). Our analysis is similar to the proof of [32, Lemma 3.7]. By the triangle inequality we have
$\|\widehat{\bm{P}}f(\bm{\Lambda})(\bm{I}-\widehat{\bm{P}})\|_{\textsf{{F}}}\leq\left\|\begin{bmatrix}\bm{0}&\\ &f(\bm{\Lambda}_{n\setminus k})\end{bmatrix}\widehat{\bm{P}}\right\|_{\textsf{{F}}}+\left\|(\bm{I}-\widehat{\bm{P}})\begin{bmatrix}f(\bm{\Lambda}_{k})&\\ &\bm{0}\end{bmatrix}\right\|_{\textsf{{F}}}.$
Using a similar argument as in (13) we have
$\left\|(\bm{I}-\widehat{\bm{P}})\begin{bmatrix}f(\bm{\Lambda}_{k})&\\ &\bm{0}\end{bmatrix}\right\|_{\textsf{{F}}}\leq\|p(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}\max\limits_{i=1,\ldots,k}\left|\frac{f(\lambda_{i})}{p(\lambda_{i})}\right|,$ (14)
and since $\bm{F}(\bm{I}+\bm{F}^{\textsf{{T}}}\bm{F})^{-1}\bm{F}^{\textsf{{T}}}\preceq\bm{F}\bm{F}^{\textsf{{T}}}$ we have
$\left\|\begin{bmatrix}\bm{0}&\\ &f(\bm{\Lambda}_{n\setminus k})\end{bmatrix}\widehat{\bm{P}}\right\|_{\textsf{{F}}}^{2}$ $=\operatorname{tr}\left(\begin{bmatrix}\bm{0}&\\ &f(\bm{\Lambda}_{n\setminus k})\end{bmatrix}\widehat{\bm{P}}\begin{bmatrix}\bm{0}&\\ &f(\bm{\Lambda}_{n\setminus k})\end{bmatrix}\right)$
$=\operatorname{tr}(f(\bm{\Lambda}_{n\setminus k})\bm{F}(\bm{I}+\bm{F}^{\textsf{{T}}}\bm{F})^{-1}\bm{F}^{\textsf{{T}}}f(\bm{\Lambda}_{n\setminus k}))$
$\leq\operatorname{tr}(f(\bm{\Lambda}_{n\setminus k})\bm{F}\bm{F}^{\textsf{{T}}}f(\bm{\Lambda}_{n\setminus k}))=\|f(\bm{\Lambda}_{n\setminus k})\bm{F}\|_{\textsf{{F}}}^{2}$
$\leq\|f(\bm{\Lambda}_{n\setminus k})\|_{2}^{2}\|p(\bm{\Lambda}_{k})^{-1}\|_{2}^{2}\|p(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2}$
$\leq\|p(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}\left|\frac{f(\lambda_{i})}{p(\lambda_{i})}\right|^{2}.$ (15)
Inserting the bounds (13), (14), and (15) into (12) and optimizing over $\mathbb{P}_{s-1}$ yields the desired inequality. ∎
3.3 Probabilistic bounds
With the structural bound available, we are ready to derive probabilistic bounds for $\|f(\bm{A})-\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}$ and $\|f(\bm{A})-\bm{Q}_{s}\llbracket\bm{Q}_{s}^{\textsf{{T}}}f(\bm{A})\bm{Q}_{s}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}$. Note that by Lemma 3.3 it is sufficient to derive a probabilistic bound for $\mathcal{E}_{\bm{\Omega}}(s;f)$ defined in Eq. 10.
We will bound $\mathcal{E}_{\bm{\Omega}}(s;f)$ in terms of a deterministic quantity
$\mathcal{E}(s;f)=\min\limits_{p\in\mathbb{P}_{s-1}}\left[\|p(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}\left|\frac{f(\lambda_{i})}{p(\lambda_{i})}\right|^{2}\right],$ (16)
which again quantifies how large a polynomial can be (relative to $f$) on the eigenvalues $\lambda_{1},\ldots,\lambda_{k}$ and small on the remaining eigenvalues. However, $\mathcal{E}(s;f)$ does not depend on the randomness used by the algorithm.
Lemma 3.4.
If $\bm{\Omega}$ is a standard Gaussian matrix, $\bm{\Omega}_{k}$ and $\bm{\Omega}_{n\setminus k}$ are as defined as in (2), and $\mathcal{E}_{\bm{\Omega}}(s;f)$ and $\mathcal{E}(s;f)$ are as defined in Eqs. 10 and 16, then
  1. (i)
    for any $u,t\geq 0$, with probability at least $1-e^{-(u-2)/4}-\sqrt{\pi k}\left(\frac{t}{e}\right)^{-(\ell-k+1)/2}$,
    $\mathcal{E}_{\bm{\Omega}}(s;f)\leq\frac{utk}{\ell-k+1}\mathcal{E}(s;f);$
  2. (ii)
    if $\ell-k\geq 2$ we have
    $\mathbb{E}[\mathcal{E}_{\bm{\Omega}}(s;f)]\leq\frac{k}{\ell-k-1}\mathcal{E}(s;f).$
Proof.
(i): For any polynomial $p\in\mathbb{P}_{s-1}$ by [40, Proposition 8.6] we have with probability at least $1-e^{-(u-2)/4}-\sqrt{\pi k}\left(\frac{t}{e}\right)^{-(\ell-k+1)/2}$
$\left[\|p(\bm{\Lambda}_{n\setminus k})\bm{\Omega}_{n\setminus k}\bm{\Omega}_{k}^{\dagger}\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}\left|\frac{f(\lambda_{i})}{p(\lambda_{i})}\right|^{2}\right]\leq\frac{utk}{\ell-k+1}\left[\|p(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}\left|\frac{f(\lambda_{i})}{p(\lambda_{i})}\right|^{2}\right].$
The inequality is respected if we minimize both sides over all polynomials.
(ii): This is proven in an identical fashion utilizing the expectation bound in [40, Proposition 8.6]. ∎
3.4 Error bounds for Krylov aware low-rank approximation
With Theorem 3.1, Lemma 3.3, and Lemma 3.4 we can now derive a probabilistic error bound for $\textsf{{ALG}}_{k}(s,r;f)$, the output of Algorithm 4. By an almost identical argument one can obtain a similar bound for $\textsf{{ALG}}(s,r;f)$, but we omit the details.
Theorem 3.5.
Consider $\bm{A}\in\mathbb{R}^{n\times n}$ as defined in (1) with smallest and largest eigenvalues $\lambda_{\min}$ and $\lambda_{\max}$ respectively. Then, with $\mathcal{E}(s;f)$ as defined in Eq. 16
  1. (i)
    with probability at least $1-e^{-(u-2)/4}-\sqrt{\pi r}\left(\frac{t}{e}\right)^{-(\ell-k+1)/2}$,
    $\|f(\bm{A})-\textsf{{ALG}}_{k}(s,r;f)\|_{\textsf{{F}}}\leq$ $4\sqrt{\ell s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|f(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}+$
    $\sqrt{\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+\frac{5utk}{\ell-k+1}\mathcal{E}(s;f)};$
  2. (ii)
    if $\ell-k\geq 2$ that
    $\mathbb{E}\|f(\bm{A})-\textsf{{ALG}}_{k}(s,r;f)\|_{\textsf{{F}}}\leq$ $4\sqrt{\ell s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|f(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}+$
    $\sqrt{\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+\frac{5k}{\ell-k-1}\mathcal{E}(s;f)}.$
Proof.
(i): By applying Theorem 3.1, Lemma 3.3, and Lemma 2.3 we obtain the following structural bound
$\begin{split}\|f(\bm{A})-\bm{Q}_{s}\llbracket f(\bm{T}_{q})_{1:d_{s,\ell},1:d_{s,\ell}}\rrbracket_{k}\bm{Q}_{s}^{\textsf{{T}}}\|_{\textsf{{F}}}\leq&4\sqrt{\ell s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|f(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}+\\ &\sqrt{\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+5\mathcal{E}_{\bm{\Omega}}(s;f)}.\end{split}$ (17)
Applying Lemma 3.4 yields (i) directly. (ii) follows from Lemma 3.4 and an application of Jensen’s inequality. ∎
We conclude this section by commenting on the three terms appearing in the bounds in Theorem 3.5. The dependence on $4\sqrt{\ell s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|f(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}$ captures the inherent cost of approximating a quadratic form with $f(\bm{A})$ using the Lanczos method, as in done in Line 3 of Algorithm 4. A similar term would also arise in an analysis of the Randomized SVD when matvecs are computed approximately using a Krylov subspace method. The $\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}$ term is due to the fact that our error can never be below the optimal rank $k$ approximation error. Finally, $\mathcal{E}(s;f)$ tells us that $\bm{Q}_{s}$ is a good orthonormal basis for low-rank approximation if there is a polynomial of degree at most $s-1$ that is large on the eigenvalues $\lambda_{1},\ldots,\lambda_{k}$ (which correspond to the top subspace of $f(\bm{A})$) and is small on the eigenvalues $\lambda_{k+1},\ldots,\lambda_{n}$.
As discussed in Section 3.2, one possible choice of polynomial would be an approximation to $f$, in which case $\mathcal{E}(s;f)$ would be close to $\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}$. In this case, ignoring the polynomial approximation term, we would recover a bound almost identical to the standard Randomized SVD error bound. In particular, when matvecs with $f(\bm{A})$ are implemented exactly, Randomized SVD can be shown to have expected error [21]:
$\sqrt{\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}+\frac{k}{\ell-k-1}\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}}.$
However, other possibly polynomials can also be chosen, so we might have $\mathcal{E}(s;f)\ll\|f(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}$. In the next section, we give an example involving the matrix exponential to better illustrate why this is often the case.
3.5 Example bounds for the matrix exponential
One matrix function for which it is often desirable to obtain a low-rank approximation is the matrix exponential, $\exp(\bm{A})$. This task arises in tasks ranging from network analysis [28], to quantum thermodynamics [6, 10], to solving PDEs [32]. In this section, we do a deeper dive into our main bounds for Krylov aware low-rank approximation when applied to the matrix exponential to better understand how they compare to bounds obtainable using the more naive approach of directly combining the randomized SVD with a black-box matrix-vector product approximation algorithm for $\exp(\bm{A})$. For conciseness of exposition, we focus on expectation bounds, and defer details of proofs to Appendix A.
We first demonstrates that, as discussed in the previous section, by choosing $p$ in $\mathcal{E}(s;\exp(x))$ to be a polynomial approximation to $\exp(x)$, we can recover the bounds of the randomized SVD [21, Theorem 10.5], up to small factors accounting for the fact that $\exp(x)$ is not a polynomial.
Corollary 3.6.
Consider the setting of Theorem 3.5 with $f(x)=\exp(x)$. Then, if $\ell-k\geq 2$ and $\gamma_{1,n}:=\lambda_{\max}-\lambda_{\min}=\lambda_{1}-\lambda_{n}$, and $s\geq e\gamma_{1,n}$ we have
$\mathbb{E}\|\exp(\bm{A})-\textsf{{ALG}}_{k}(s,r;\exp(x))\|_{\textsf{{F}}}$
$\hskip 20.00003pt\leq\frac{\sqrt{\ell s}\gamma_{1,n}^{2r+2}}{2^{4r+1}(2r+2)!}\|\exp(\bm{A})\|_{2}+\sqrt{1+\frac{1}{(1-\frac{\gamma_{1,n}^{s}}{s!})^{2}}\frac{5k}{\ell-k-1}}\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}.$
Note that as $\frac{\sqrt{\ell s}\gamma_{1,n}^{2r+2}}{2^{4r+1}(2r+2)!}\|\exp(\bm{A})\|_{2}\to 0$ and $\frac{\gamma_{1,n}^{s}}{s!}\to 0$, we recover the classical bound for the randomized SVD [21, Theorem 10.5]. Furthermore, these two terms converge superexponentially to 00 as $s,r\to+\infty$. However, due to the $\frac{5k}{\ell-k-1}$, the above bound only implies a subexponential convergence to the error of the optimal rank $k$ approximation error, $\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}$, in terms of the number of required matrix-vector products with $\bm{A}$, which scales as $\ell(r+s)$.
However, as argued in Section 3.2 there are other polynomials that can achieve a significantly tighter upper bound of $\mathcal{E}(s;\exp(x))$. For example, by choosing $p$ in $\mathcal{E}(s;\exp(x))$ to be a scaled and shifted Chebyshev polynomial we can show the output of Algorithm 4 converges exponentially in $s$ to the optimal low-rank approximation of $\exp(\bm{A})$, even if $\ell$ stays fixed. This demonstrates why Algorithm 4 is expected to return a better low-rank approximation compared to applying the randomized SVD immediately on $\exp(\bm{A})$. In particular, we have the following result.
Corollary 3.7.
Consider the setting of Theorem 3.5 with $f(x)=\exp(x)$. Define $\gamma_{i,j}=\lambda_{i}-\lambda_{j}$ to be the gap between the $i^{\text{th}}$ and $j^{\text{th}}$ largest eigenvalues. Then, if $\ell-k\geq 2$ and $s\geq 2+\frac{\gamma_{1,k}}{\log\left(\frac{\gamma_{1,n}}{\gamma_{k,n}}\right)}$
$\mathbb{E}\|\exp(\bm{A})-\textsf{{ALG}}_{k}(s,r;\exp(x))\|_{\textsf{{F}}}\leq\frac{\sqrt{\ell s}\gamma_{1,n}^{2r+2}}{2^{4r+1}(2r+2)!}\|\exp(\bm{A})\|_{2}$
$\hskip 50.00008pt+\sqrt{1+\exp(2\gamma_{k,n})\cdot 2^{-2(s-2)\sqrt{\min\left\{1,2\frac{\gamma_{k,k+1}}{\gamma_{k+1,n}}\right\}}}\frac{80k}{\ell-k-1}}\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}.$
4 Numerical experiments
In this section we compare the Krylov aware low-rank approximation (Algorithm 4) and Algorithm 2 (assuming exact matvecs with $f(\bm{A})$) and Algorithm 3 (inexact matvecs with $f(\bm{A})$). All experiments have been performed in MATLAB (version 2020a) and scripts to reproduce the figures are available at https://github.com/davpersson/Krylov_aware_LRA.git.
4.1 Test matrices
We begin with outlining the test matrices and matrix functions used in our examples.
4.1.1 Exponential integrator
The following example is taken from [32]. Consider the following parabolic differential equation
$\begin{split}&u_{t}=\kappa\Delta u+\lambda u\text{ in }[0,1]^{2}\times[0,2]\\ &u(\cdot,0)=\theta\text{ in }[0,1]^{2}\\ &u=0\text{ on }\Gamma_{1}\\ &\frac{\partial u}{\partial\bm{n}}=0\text{ on }\Gamma_{2}\end{split}$
for $\kappa,\lambda>0$ and $\Gamma_{2}=\{(x,1)\in\mathbb{R}^{2}:x\in[0,1]\}$ and $\Gamma_{1}=\partial\mathcal{D}\setminus\Gamma_{2}$. By discretizing in space using finite differences on a $100\times 100$ grid we obtain an ordinary differential equation of the form
$\begin{split}\dot{\bm{u}}(t)&=\bm{A}\bm{u}(t)\text{ for }t\geq 0,\\ \bm{u}(0)&=\bm{\theta},\end{split}$ (18)
for symmetric matrix $\bm{A}\in\mathbb{R}^{9900\times 9900}$. It is well known that the solution to (18) is given by $\bm{u}(t)=\exp(t\bm{A})\bm{\theta}$. Suppose that we want to compute the solution for $t\geq 1$. One can verify that
$\max\limits_{t\geq 1}\frac{\|\exp(t\bm{A})-\llbracket\exp(t\bm{A})\rrbracket_{k}\|_{\textsf{{F}}}}{\|\exp(t\bm{A})\|_{\textsf{{F}}}}=\frac{\|\exp(\bm{A})-\llbracket\exp(\bm{A})\rrbracket_{k}\|_{\textsf{{F}}}}{\|\exp(\bm{A})\|_{\textsf{{F}}}},$
and it turns out that $\exp(\bm{A})$ admits a good rank $60$ approximation
$\frac{\|\exp(\bm{A})-\llbracket\exp(\bm{A})\rrbracket_{60}\|_{\textsf{{F}}}}{\|\exp(\bm{A})\|_{\textsf{{F}}}}\approx 4\times 10^{-4}.$
Hence, we can use Algorithm 4 to compute $\bm{Q}_{s}$ and $\bm{T}_{q}$ and use them to efficiently construct a rank $60$ approximation to $\exp(t\bm{A})$ for any $t$ with almost no additional cost.
In the experiments we set $\kappa=0.01$ and $\lambda=1$.
4.1.2 Estrada index
For an (undirected) graph with adjacency matrix $\bm{A}$ the Estrada index is defined as $\operatorname{tr}(\exp(\bm{A}))$. It is used to measure the degree of protein folding [12]. One can estimate the Estrada index of a network by the Hutch++ algorithm or its variations [6, 10, 28, 31], which requires computing a low-rank approximation of $\exp(\bm{A})$. Motivated by the numerical experiments in [28] we let $\bm{A}$ be the adjacency matrix of Roget’s Thesaurus semantic graph [1].
(a) Example from Section 4.1.1
(b) Example from Section 4.1.2
(c) Example from Section 4.1.3
(d) Example from Section 4.1.4
Figure 1: Comparing relative error (19) for the the approximations returned by Algorithm 4 without truncation (untruncated), Algorithm 4 with truncation back to rank $k$ (truncated), Algorithm 3, and Algorithm 2. The black line shows the optimal rank $k$ approximation relative Frobenius norm error. The rank parameter $k$ is visible as titles in the figures. In all experiments we set $\ell=k$.
4.1.3 Quantum spin system
We use an example from [10, Section 4.3], a similar example is found in [6], in which we want to approximate $\exp(-\beta\bm{A})$ where
$\bm{A}=-\sum\limits_{i=1}^{N-1}\bm{Z}_{i}\bm{Z}_{i+1}-h\sum\limits_{i=1}^{N}\bm{X}_{i}\in\mathbb{R}^{n\times n},$
where
$\bm{X}_{i}=\bm{I}_{2^{i-1}}\otimes\bm{X}\otimes\bm{I}_{2^{N-i}},\quad\bm{Z}_{i}=\bm{I}_{2^{i-1}}\otimes\bm{Z}\otimes\bm{I}_{2^{N-i}}$
where $\bm{X}$ and $\bm{Z}$ are the Pauli operators
$\bm{X}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\quad\bm{Z}=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}.$
Estimating the partition function $Z(\beta)=\operatorname{tr}(\exp(-\beta\bm{A}))$ is an important task in quantum mechanics [35], which once again can benefit from computing a low-rank approximation of $\exp(-\beta\bm{A})$.
In the experiments we set $N=14$ so that $n=2^{14}$, $\beta=0.3$, and $h=10$.
4.1.4 Synthetic example for the matrix logarithm
We generate a symmetric matrix $\bm{A}\in\mathbb{R}^{5000\times 5000}$ with eigenvalues $\lambda_{i}=\exp(\frac{1}{i^{2}})$ for $i=1,\ldots,n$. We let $f(x)=\log(x)$ so that the eigenvalues of $f(\bm{A})$ are $f(\lambda_{i})=\frac{1}{i^{2}}$ for $i=1,\ldots,n$.
(a) Example from Section 4.1.1
(b) Example from Section 4.1.2
(c) Example from Section 4.1.3
(d) Example from Section 4.1.4
Figure 2: Comparing relative error (19) for the the approximations returned by Algorithm 4 without truncation (untruncated), Algorithm 4 with truncation back to rank $k$ (truncated), Algorithm 3, and Algorithm 2. The black line shows the optimal rank $k$ approximation relative Frobenius norm error. The rank parameter $k$ is visible as titles in the figures. In all experiments we set $\ell=k+5$.
4.2 Comparing relative errors
In this section we compare the error returned by Algorithm 4, Algorithm 2, and Algorithm 3. If $\bm{C}$ is a low-rank approximation returned by one of the algorithms then we compare the relative error
$\frac{\|f(\bm{A})-\bm{C}\|_{\textsf{{F}}}}{\|f(\bm{A})\|_{\textsf{{F}}}}.$ (19)
In all experiments we set the parameters in Algorithm 4 and Algorithm 3 to be $\ell=k$ or $\ell=k+5$ and $s=r$ so that the total number of matrix vector products with $\bm{A}$ is $2\ell s$. When we run Algorithm 3 we compute matvecs with $f(\bm{A})$ exactly, which cannot be done in practice. Hence, the results from this algorithm are only used as a reference for Algorithm 4 and Algorithm 3. The results are presented in Figure 1 for $\ell=k$ and Figure 2 for $\ell=k+5$. All results confirm that Algorithm 4 returns a more accurate approximation than Algorithm 3, and can even be more accurate than the Algorithm 2. Note that for the example given in Section 4.1.3 the error for the untruncated version of the approximation returned by Algorithm 4 stagnates. This is because the error from the approximation of the quadratic form dominates the error. In this case, $r$ should be chosen larger than $s$. However, we leave it as an open research question to determine the optimal balancing of $s$ and $r$.
5 Conclusion
In this work we have provided an analysis of the Krylov-aware low-rank approximation presented in [6]. In particular, we have shown that the Krylov-aware low-rank approximation to $f(\bm{A})$ is never worse than the approximation returned by the randomized SVD on $f(\bm{A})$, up to a polynomial approximation error factor. Furthermore, the numerical experiments demonstrate that the Krylov-aware low-rank approximation is significantly more accurate than the approximation returned by the randomized SVD on $f(\bm{A})$.
Acknowledgements
This work was done while David Persson was a PhD-student at EPFL and supported by the SNSF research project Fast algorithms from low-rank updates, grant number: 200020_178806. This work is also contained in his PhD-thesis [34]. Tyler Chen and Christopher Musco were partially supported by NSF Award #2045590. The authors thank Daniel Kressner for helpful comments on this work.
Appendix A Proofs of Corollary 3.6 and Corollary 3.7
Proof of Corollary 3.6.
First note that by a standard Chebyshev interpolation bound [39, Lecture 20]
$\inf\limits_{p\in\mathbb{P}_{2r+1}}\|\exp(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}\leq\frac{\gamma_{1,n}^{2r+r}}{2^{4r+3}(2r+2)!}\|\exp(\bm{A})\|_{2}.$
Hence,
$4\sqrt{\ell s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|f(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}\leq 4\sqrt{\ell s}\frac{\gamma_{1,n}^{2r+2}}{(2r+2)!}\|\exp(\bm{A})\|_{2}.$
We proceed with bounding $\mathcal{E}(s;\exp(x))$. Let $p(x)=\sum\limits_{i=1}^{s-1}\frac{(x-\lambda_{\min})^{i}}{i!}$. Then for $x\in[\lambda_{\min},\lambda_{\max}]$ we have $0\leq p(x)\leq\exp(x-\lambda_{\min})$. Consequently, $\|p(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}\leq\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}\exp(-\lambda_{\min})$. Furthermore, for $x\in[\lambda_{\min},\lambda_{\max}]$ we have
$0\leq\exp(x-\lambda_{\min})-p(x)\leq\frac{(x-\lambda_{\min})^{s}}{s!}\exp(x-\lambda_{\min}).$
Note that by the assumption on $s$ and a Stirling approximation $s!\geq\sqrt{2\pi s}\left(\frac{s}{e}\right)^{s}$ [37] we have $\frac{\gamma_{1,n}^{s}}{s!}<1$. Hence,
$\max\limits_{i=1,\ldots,k}\left|\frac{\exp(\lambda_{i})}{p(\lambda_{i})}\right|$ $=\exp(\lambda_{\min})\max\limits_{i=1,\ldots,k}\left|\frac{\exp(\lambda_{i}-\lambda_{\min})}{p(\lambda_{i})}\right|$
$=\exp(\lambda_{\min})\left(1-\frac{(\lambda_{\max}-\lambda_{\min})^{s}}{s!}\right)^{-1}.$ (20)
Therefore, $\mathcal{E}(s;\exp(x))$ is bounded above by
$\mathcal{E}(s;\exp(x))\leq\frac{1}{(1-\frac{(\lambda_{\max}-\lambda_{\min})^{s}}{s!})^{2}}\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}=\frac{1}{(1-\frac{\gamma_{1,n}^{s}}{s!})^{2}}\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}.$ (21)
Plugging the inequalities (A) and (21) into Theorem 3.5 yields the desired inequality. ∎
Proof of Corollary 3.7.
Bounding $4\sqrt{\ell s}\inf\limits_{p\in\mathbb{P}_{2r+1}}\|\exp(x)-p(x)\|_{L^{\infty}([\lambda_{\min},\lambda_{\max}])}$ is done identically to in Corollary 3.6. We proceed with bounding $\mathcal{E}(s;\exp(x))$. Define $\delta(x)=\frac{x-\lambda_{k+1}}{\lambda_{k+1}-\lambda_{n}}$. Define the polynomial $p(x)=(1+x-\lambda_{n})T_{s-2}\left(1+2\delta(x)\right)$, where $T_{s-2}$ is the Chebyshev polynomial of degree $s-2$. Hence, recalling the definition Eq. 16 of $\mathcal{E}(s;\exp(x))$, since $0\leq 1+x\leq\exp(x)$ for $x\geq 0$ and $|T_{s-2}(x)|\leq 1$ for $x\in[-1,1]$ we have,
$\|p(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}\leq\|\exp(\bm{\Lambda}_{n\setminus k}-\lambda_{n}\bm{I})\|_{\textsf{{F}}}$
Hence, using that $\frac{\exp(x)}{1+x}\leq\exp(x)$ for $x\geq 0$ we get
$\mathcal{E}(s;\exp(x))$ $\leq\|p(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}\left|\frac{\exp(\lambda_{i})}{p(\lambda_{i})}\right|^{2}$
$\leq\|\exp(\bm{\Lambda}_{n\setminus k}-\lambda_{n}\bm{I})\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}e^{2\lambda_{n}}\left|\frac{\exp(\lambda_{i}-\lambda_{n})}{p(\lambda_{i})}\right|^{2}$
$=\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}\left|\frac{\exp(\lambda_{i}-\lambda_{n})}{p(\lambda_{i})}\right|^{2}$
$\leq\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}\max\limits_{i=1,\ldots,k}\left|\frac{\exp(\lambda_{i}-\lambda_{n})}{T_{s-2}\left(1+2\delta(\lambda_{i})\right)}\right|^{2}.$ (22)
First note that for $|x|\geq 1$ we have
$T_{s-2}(x)=\frac{1}{2}\left(\left(x+\sqrt{x^{2}-1}\right)^{s-2}+\left(x-\sqrt{x^{2}-1}\right)^{s-2}\right).$
Hence, for $\delta\geq 0$ we have $|T_{s-2}(1+2\delta)|\geq\frac{1}{2}(1+2\delta+2\sqrt{\delta+\delta^{2}})^{s-2}=:\frac{1}{2}h(\delta)^{s-2}$. Therefore,
$\max\limits_{i=1,\ldots,k}\left|\frac{e^{\lambda_{i}-\lambda_{n}}}{T_{s-2}\left(1+2\delta(\lambda_{i})\right)}\right|^{2}$ $\leq 4\max\limits_{i=1,\ldots,k}\left|\frac{\exp(\lambda_{i}-\lambda_{n})}{h(\delta(\lambda_{i}))^{s-2}}\right|^{2}$
$=4\max\limits_{i=1,\ldots,k}\left(\exp(\lambda_{i}-\lambda_{n}-(s-2)\log(h(\delta(\lambda_{i}))))\right)^{2}.$
Note that the function
$g(x):=x-\lambda_{n}-(s-2)\log(h(\delta(x)))$
is convex in $[\lambda_{k},\lambda_{1}]$ because $\log(h(x))$ is concave for $x\geq 0$ and $\delta(x)$ is linear positive function for any $x\geq\lambda_{k}$. Hence, the maximum of $g$ on the interval $[\lambda_{k},\lambda_{1}]$ is attained at either $x=\lambda_{k}$ or $x=\lambda_{1}$. The maximum of $g$ is attained at $x=\lambda_{k}$ whenever $g(\lambda_{1})\leq g(\lambda_{k})$ which happens whenever
$\frac{\lambda_{1}-\lambda_{k}}{\log\left(\frac{h(\delta(\lambda_{1}))}{h(\delta(\lambda_{k}))}\right)}+2\leq s.$ (23)
Note that we have whenever $x\geq y\geq 0$ we have $\frac{h(x)}{h(y)}\geq\frac{1+x}{1+y}$, which can be seen by noting that $\frac{h(x)}{1+x}$ is an increasing function. Hence, (23) holds whenever
$\frac{\lambda_{1}-\lambda_{k}}{\log\left(\frac{\lambda_{1}-\lambda_{n}}{\lambda_{k}-\lambda_{n}}\right)}+2\leq s,$
which is our assumption on $s$. Hence,
$\max\limits_{i=1,\ldots,k}\left|\frac{\exp(\lambda_{i}-\lambda_{n})}{T_{s-2}\left(1+2\delta(\lambda_{i})\right)}\right|^{2}\leq 4\frac{\exp(2(\lambda_{k}-\lambda_{n}))}{h(\delta(\lambda_{k}))^{2(s-2)}}.$ (24)
Now we use the argument from [29, p.21], which shows that $h(\delta)^{(s-2)}\geq 2^{\sqrt{\min\{1,2\delta\}}(s-2)-1}$. Plugging this inequality into Eq. 24 and then Eq. 22 and using the definition for $\gamma_{i,j}$ yields
$\mathcal{E}(s;\exp(x))\leq 16\exp(2\gamma_{k,n})2^{-2(s-2)\sqrt{\min\left\{1,2\frac{\gamma_{k,k+1}}{\gamma_{k+1,n}}\right\}}}\|\exp(\bm{\Lambda}_{n\setminus k})\|_{\textsf{{F}}}^{2}.$